The equation W=120I-6I^2 represents the power (W) in watts that a 120 volt circuit can deliver if it has a resistance of 6 ohmmeter. What is the maximum power, in watts, that can be delivered by this circuit?
My name is Chantal and I'm not sure how to do this. I was sick when he taught us word problems. I believe the answer is the vertex seeing that it wants a maximum but I'm not sure of the equation. Is it w= -6I^2 + 120 + 6 or is it
6=-6I^2 + 120 + 0.
If you can help me it would be greatly appreciated. Please and Thank you.This is not from a text book.
Chantal

Answer by Alan3354(69443)

The equation W=120I-6I^2 represents the power (W) in watts that a 120 volt circuit can deliver if it has a resistance of 6 ohms. What is the maximum power, in watts, that can be delivered by this circuit?
My name is Chantal and I'm not sure how to do this. I was sick when he taught us word problems. I believe the answer is the vertex seeing that it wants a maximum but I'm not sure of the equation. Is it w= -6I^2 + 120 + 6 or is it
6=-6I^2 + 120 + 0.
--------------------
Set 120I - 6I^2 = 0 and find the vertex
The max is at I = 10 amps.
W = 1200 - 600 = 600 watts